OpenStudy (anonymous):

How would I go about writing the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form ?

OpenStudy (anonymous):

First, consider the slope of the line y=4x+7. What will be the slope of a line perpendicular to it?

OpenStudy (anonymous):

slope of y = 4x + 7 is 4 implies slope of desired line is -1/4 (m1* m2 = -1) so the eqn will be y+3 = -1/4(x-2) 4y + 12= 2-x 4y = -x - 10 y = -1/4(x +10)

OpenStudy (anonymous):

or y = -1/4x -5/2

OpenStudy (anonymous):

Tell her why at least, Tapan. If two lines are perpendicular, then their slopes will be the opposite reciprocal. Reciprocal means flip it, opposite means switch the sign.

OpenStudy (anonymous):

thanks you both so much. What about the equation of the line which passes through (2, 1) and is perpendicular to x = –2 ? it doesnt ask for a certain form though..

OpenStudy (anonymous):

x = -2 has a slope equal to 90 degrees so d slope of perp line will be = 0 that is parallel to x- axis so eqn will be y=1